OpenStudy (anonymous):

Let y=e^(x/9). a) Find the differential dy. b) Evaluate dy if x=0 and dx=−0.03.

hartnn (hartnn):

do you know chain rule ? you'll need that for part a)

OpenStudy (anonymous):

Yeah I got the derivative e^(x/9)(1/9x-1) is this right?

OpenStudy (anonymous):

I can't find online calculators for differentiation to check my work

hartnn (hartnn):

how you got (1/9x -1) part ?

OpenStudy (anonymous):

Sorry (1/9)x^(1/9)-1))

hartnn (hartnn):

let me show you the chain rule thingy, \(\large [e^{f(x)}]'=e^{f(x)}f'(x)\) because derivative of e^x is just e^x got that ? here f(x) is just x/9

hartnn (hartnn):

so you would just multiply the derivative of (x/9) (which is...?) to e^(x/9)

hartnn (hartnn):

(1/9)x^(1/9)-1))<-----from this i can see that you have done the 1/9 part correct, but you have used x^n formula, which is not the function here, we need e^x formula here..

OpenStudy (anonymous):

oh ok so y' would equal (x/9)e^(x/9) because I use the power rule on the e, and e doesn't really change in its derivative beside pulling down the fraction?

hartnn (hartnn):

the derivative of x/9 is just 1/9 so, you'll get \(dy=\large \dfrac{1}{9}e^{x/9}dx\) got this ?

OpenStudy (anonymous):

Yes so this is my total derivative of y=e^(x/9)?

OpenStudy (anonymous):

And I would just plug in the numbers of part b into my derivative?

hartnn (hartnn):

yes..,more precisely, its dy and yes, you just plug in x=0 and dx =-0.03 here.

hartnn (hartnn):

ask if any doubts anywhere :)

OpenStudy (anonymous):

I put (1/9)e^(x/9) as my answer and the assignment tells me it's still wrong. It's an online assignment.

hartnn (hartnn):

you missed out dx (1/9)e^(x/9) dx

OpenStudy (anonymous):

oh ok I got the problem now thank you so much

hartnn (hartnn):

welcome ^_^